OpenStudy (anonymous):

6+3t=8t-14 (please show work)

OpenStudy (anonymous):

3t-8t=-14-6 .....-5t=-20 ......t=4

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

6 + 3t = 8t - 14 -6 -6 3t = 8t - 20 -8t -8t -5t = - 20 divide -20 to both sides t=4